OpenStudy (anonymous):

Solve for x using the Power Property of Logarithms. 7^(3+x)=54 A)x≈ 2.95 B)x≈ 0.95 C)x≈ -0.95 D)x≈ -2.95

OpenStudy (anonymous):

anybody ? D: im strugglimg so much with this !

OpenStudy (ybarrap):

Take logs on both sides: $$ 7^{3+x}=54\\ \ln7^{3+x}=\ln54\\ (3+x)\ln7=\ln54\\ 3+x=\cfrac{\ln54}{\ln7}\\ x=\cfrac{\ln54}{\ln7}-3\\ $$ Does this make sense?